OpenStudy (anonymous):

let g(x)= x+3/2x-5 find g(-3) Can the denominator be a negative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it just can't be zero

OpenStudy (anonymous):

would my answer be -(1/11) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[g(x)=\frac{x+3}{2x-5}\] \[g(-3)=\frac{-3+3}{2\times (-3)-5}=\frac{0}{-11}=0\]

OpenStudy (anonymous):

but that doesn't mean its undefined?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{a}{0}\] is undefined but \(\frac{0}{a}\) is \(0\)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks